A source emitting a frequency of 480 Hz is moving toward a stationary observer. If the frequency heard by the observer is 560 Hz, what is the speed of the source? Assume a sound speed of 345 m/s.



1
Expert's answer
2021-07-28T09:21:07-0400

The emitted "f_0 = 480Hz" and received "f=560Hz" frequencies are connected with the following formula (see Doppler effect):


"f = f_0\dfrac{1}{1-\dfrac{v}{c}}"

where "c = 545m\/s" is the sound speed, and "v" is the speed of the source with respect to the medium. Thus, expressing "v", obtain:

"v = c\left(\dfrac{1-f_0}{f} \right)\\
v = 545m/s\left(1- \dfrac{480Hz}{560Hz}\right) \approx 77.9m/s"

Answer. 77.9 m/s.